3x - y = -3
2x + y = -7
3x = 10 - 2y
2x = 5 + 2y



Answer by Alan3354(69443)

3x - y = -3
2x + y = -7
------------- Add
5x = -10
x = -2
-------
Sub x into eqn 1
-6 - y = -3
y = -3
---------
If these 2 are another problem, do it the same way.
3x = 10 - 2y
2x = 5 + 2y